Please proof it
Pros Usines pls):l. P(AUB)- p(A)+ p(B)
Answer #1

Ans:

As,

A U B = (A n B) U (A n B’) U (A’ n B)

So,

P(A U B)=P(A n B) + P(A n B’) + P (A’ n B) +P (A n B) - P (A n B)

Now,

P(A n B) + P(A n B’)=P(A)

P (A’ n B) +P (A n B)=P(B)

Hence,

P(A U B)= P(A)+P(B)-P(A n B)